OpenStudy (blackops2luvr):

7(5 + 9) = 7 · x + 7 · 9

OpenStudy (anonymous):

First work in all the parenthesis. Since it's one, we're going to add 4 and 9, which is 13. Now it's 7(13) = 7*x+7*9 Now, 7( can be written as 7*( since we always multiply parenthesis if there's no + or - next to it. So let's go ahead and multiply everything out. 7*13 = 91 7*x = 7x (since they can't combine) 7*9 = 63 Now plug in everything back. 91 = 7x + 63 Subtract both sides by 63 to get 7X by itself 91 - 63 = 7x 28 = 7x Divide both sides by 7 4 = x We can flip it over x = 4

OpenStudy (blackops2luvr):

its 5

OpenStudy (blackops2luvr):

A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (loser66):

to me, I will work from the right hand side, factor 7 out , I have 7 (x +9) now , compare to the left hand side 7 (5 + 9 ) = 7 (x + 9) , easily to see x = 5

OpenStudy (anonymous):

My bad, I added 9 + 5* wrong because I thought it was a 4, not a 5 at the beginning. But you can basically use my explanation if it's required and swap out the 13 with 14.